OpenStudy (anonymous):

Laplace Transform of (1-e^-t)u(t-3)

OpenStudy (anonymous):

\[ (1-e^{-t})u(t-3)\]

OpenStudy (anonymous):

@Loser66 Lol, I was going to ask you.

OpenStudy (loser66):

I forgot all the stuff. ha!!

OpenStudy (loser66):

woah... you post the whole book ??

OpenStudy (anonymous):

How are you in high school!

OpenStudy (accessdenied):

I'm not remembering the specific identity with multiplying by the heaviside step function, but by definition u(t - 3) is neutralizing any values below t < 3. So the Laplace transform of (1 - e^(-t)) u(t - 3) is equal to: \( \displaystyle \int_{0}^{\infty} (1 - e^{-t} ) u(t - 3) e^{-st} \ dt \) = \( \displaystyle \int_{\color{blue}{3}}^{\infty} (1 - e^{-t} ) e^{-st} \ dt \) We would need to substitute u = t - 3 so that the adjustment brings back our lower boundary of 0, which is the definition of the Laplace transform of our function after substitution